Shift Summary 12 June 1999, evening.

Nickolas Solomey (solomey@hep.uchicago.edu)
Sun, 13 Jun 1999 00:00:21 -0500

Shift Summary 12 June 1999

On shift: Albert Erwin, Nick Solomey

Others present: Rick K., Rick C., Ed. B., Julie W.

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 3hrs debugging csi
channel

Accounting of lost time: No beam from Tevatron, Tevatron studies.

========================================================================
Runs ended this shift: Several for debugging CsI bad channel, but nothing
else worth mentioning.

------------------------------------------------------------------------
Tapes brought to FCC this shift: None

========================================================================
Accesses during shift: 17:15-23:35 Ed and Nick made access to swap bad
1443 HV module in CsI array.

========================================================================
Overview of shift: No beam so we did some testing of system and accesses
for CSI.

========================================================================
Description of Shift:
15:30> Arrived on shift, found that TeVatron studies were underway so no
beam to KTeV, but expect to get beam back somewhere between 16:00 to 18:00
hrs.

16:15> RC called, said that tevatron would be an hour still so when
it comes back up page him.

16:20> RF called, saying that after we page RC page him also.

17:25> MCR called saying no beam until 7pm., RC was disapointed.

18:00> Started run even though there was no beam to make sure everything
was still working, found that two patches of 8 channels that have an 8%
shift in gain, Ed investigating. DC4 illumination seems to be missing, but
that might be just low statistics.

18:30> Ed reports after he did a reinitilization of pipeline that fixed
those two patches of 8 gain shifted channels.

18:45> Those two patches of 8 bad channels in CsI array persist, Ed
does not understand so he paged Peter, who keeps calling on his cellphone
and gets cut off.

19:00> Problem with those 16 channels tracked with Peter's help by phone
to a 1443 HV module, these voltages look messed up so we will try to
reload, if not then we have to swap the HV module out with a better one.

19:35> Ed and Nick make access, swap bad 1443 HV module, after turning
triggers back on the 16 bad channels are now good.

20:15> CsI channel 2076 drifting by 46%, HV for that channel does
not verify, tried resetting HV from DB and HV comes up but then goes
off again. Talked with Peter who said he has no special help, call
CsI on call. We paged Julie. This is one of the channels in the HV
card that was swapped.

21:15> Julie and Nick make access, change the HV module again. This
has 2 channels with >5% gain drift but the two groups of 8 channels
are back like when Ed first notied it at 18:00 hrs. Julie investagating.

21:20> Problem may be a bad slot in main frame of 1440. Return downstairs
to cable swap, etc.

21:50> Rick Coleman calls. There has been a quench. Stuck valve is
usually an hour problem. He is going home. Rick Ford might try to
tune from home if they come up again by 12 or 1 am. MCR operators
are tired too. They may restore last night's values and return beam
to us if they give up for the day.

22:15> Elizabeth T. changed REG HV's after another calibration pass.

22:30> Rick Ford says there will be no beam until midnight, so he
will not do any beam tuning tonight.

22:40> Julie and Ed make an access the bad HV slot may be due to
a bad divider on the PMT so they will swap it, maybe (this is a
wild guess) but it is possible that a bad divider is killing that
channel in the HV 1443 which then effect the whole card?

23:07> Ed and Julie out of access, swapped the divider now they
check the CsI array for this persistent problem. Channel 2076 looks
fixed, but two others far away from this area now need to be
watched 1247 and 2753.

23:30> CsI channel 1247 recovered and seems fine, while channel
2753 is recovering more slowly and hope that it will be alright
in an hour or so.

23:31> DEMACs terminal the PC that runs the blockhouse temp. controls
etc. crashed. Needs rebooting by its master.

23:50> CsI channel 2877 just went 50% out of gain, may be a bad
divider that they gave us last week when this one was swapped out.
We need to watch this one over the next few 10-20 min according to
Ed on the phone, more news on the next shift.

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey